# Tripartite entanglement and Bell non-locality in loop-induced Higgs boson decays

### v2

@@ 1 Introduction
In this work, we are interested on entanglement between the helicity degrees of freedom of the particles in the tripartite final state corresponding to the Higgs boson decaying into a photon and a lepton-pair. Previously, entanglement properties of bipartite Higgs decays were analyzed for the 2-qubit final states [13, 14, 15], corresponding to tau-lepton pair and two photons. Higgs boson decaying into massive gauge bosons, i.e. 2-qutrits systems, were studied in [16, 17, 18, 19, 20, 21, 22]. Also, 2-qutrits systems were explored in diboson production at LHC [20, 19, 23, 24] and through vector boson scattering [25].
On the other hand, tripartite entanglement within the HEP context is in development: Ref. [26] explores positronium decaying into three photons, QED scattering processes with spectator particle were analyzed in [27, 28, 29], Ref. [30] studies heavy fermion decaying into three fermions via generic (pseudo)scalar, (pseudo)vector and (pseudo)tensor interactions, Ref. [31] explores three-flavor entanglement in neutrino oscillations, and Ref. [32] addresses entanglement among the two spins and total angular momentum in $H\to ZZ,WW$. General properties of multipartite systems were presented in [33, 34]. Furthermore, the tripartite Higgs boson decays considered here can be related to bipartite ones, in order to investigate post-decay entanglement [35] and autodistillation phenomena [36, 37].
@@ 2 Three-body Higgs boson decays
The last sum involves the 8 helicity amplitudes (see Appendix B). The polar angle $\theta_{\gamma l}$ belongs to $[0,\pi]$ and the Mandelstam variable $s$ has relevant limits
@@
[…]
@@
$|\psi\rangle$ $\simeq$ $\frac{-i}{N}\left(8A_{\rm Tree}\kappa_{\rm CP}^{l}e^{-i\delta_{\rm CP}^{l}}\frac{s}{(m_{\rm H}^{2}-s)\sin(\theta_{\gamma l})}|+++\rangle+a_{2}(m_{\rm H}^{2}-s)\sqrt{s}(1+\cos(\theta_{\gamma l}))|++-\rangle\right.$ (3.5)
@@
×$\hskip 34.1433pt\times\left(c_{\gamma l}^{2}\left(4m_{l}^{2}-s\right)\left(s\left(s-16m_{l}^{2}c_{\rm CP}^{2}\right)+m_{\rm H}^{4}+2m_{\rm H}^{2}s\right)\right.$
@@
Concerning the one-to-one concurrences, they require a 4$\times$4 diagonalization and analytical expressions result just by neglecting lepton mass terms. In that regime, both photon-to-lepton $\mathcal{C}_{\gamma l}$ and photon-to-antilepton $\mathcal{C}_{\gamma\bar{l}}$ exactly vanish for both tree level and 1-loop contributions (then lepton mass terms in the spinors will be relevant in the numerical analysis) and the lepton-to-antilepton $\mathcal{C}_{l\bar{l}}$ for the tree level is very compact
@@
The previous two classes of entanglement (one-to-other and one-to-one) are inequivalent. In fact, the $i$-to-$jk$ entanglement limits the entanglements $i$-$j$ and $i$-$k$ by means of the Coffman-Kundu-Wootters (CKW) monogamy inequality [66]
@@ 4.2 $H\to\gamma Z$ and post-decay entanglement
where $\{k^{\mu},s_{1}\}$ and $\{q^{\nu},\bar{s}_{1}\}$ are the momentum and helicity of the photon and $Z$ boson, $\varepsilon$’s are their polarization vectors and the 3-legs form factor $V_{H\gamma Z}$ accounts for the 1-loop contribution (that only depends on $k$ and $q$ momenta). The precise dependence of this form factor on the momenta is not relevant for the following discussion. In the {+,}$\{+,-\}\otimes\{+,0,-\}$ basis, the resulting density matrix is
@@
which is independent of the form factor (since it is a global factor of the helicity amplitudes). The concurrence for a bipartite system with arbitrary dimension $d_{1}\otimes d_{2}$ was defined in [80] and the theoretical maximum is $\sqrt{2(d-1)/d}$ with $d={\rm min}\{d_{1},d_{2}\}$. For this bipartite decay, the concurrence achieves the maximum 1 and the $\gamma Z$ state (coming from the Higgs boson) results maximally entangled. However, the 2$\otimes$3 generalized CHSH operator [81, 16, 25] for this decay never exceeds 2.
@@ 5 Numerical Results
In this section, we present the distributions of the entanglement and Bell non-locality measures in the $[m_{l\bar{l}},\cos(\theta_{\gamma l})]$ plane of the phase-space. We analyze the three lepton families separately since the 1-loop contribution dominates different energy regimes in each case, and also corresponds to separate experimental channels. The numerical results correspond to the SM, i.e. $\kappa_{\rm CP}^{l}=1$ and $\delta_{\rm CP}^{l}=0$ in Eq. (B.1), including lepton mass effects and using the corresponding form factors for each kind of computation in Eq. (2.4). Then, in the plots, the range for $\cos(\theta_{\gamma l})$ is $[0,1]$ since the distributions result symmetric under $\cos\to-\cos$, and the dilepton invariant mass covers the range in Eq. (2.6).
@@
Figure 3: Entanglement and non-locality quantifiers for $H\to\gamma\tau\bar{\tau}$ in the $[m_{\tau\bar{\tau}},\cos(\theta_{\gamma\tau})]$ plane, corresponding to the full computation (first row) and two-body intermediate Higgs decay (second row). Dashed lines are contours corresponding to 0.1 and 0.9 for concurrences and to 2 and 3.5 for Mermin operator.
@@ 5.1 Tau-lepton case
The $H\to\gamma\tau\bar{\tau}$ process is dominated by the 1-loop (tree level) contribution for energies below (above) 30 GeV, as can be seen from the first row of Fig. 2. Around the $Z$-pole, both contributions interfere constructively but the resonance comes from the 1-loop diagrams type (f). The resulting one-to-other and one-to-one concurrences $\mathcal{C}_{l\bar{l}|\gamma}$ and $\mathcal{C}_{l\bar{l}}$ in the full computation are shown in the first row of Fig. 3, together with the Mermin operator $\mathcal{M}_{3}$ in the $[m_{\tau\bar{\tau}},\cos(\theta_{\gamma\tau})]$ plane.
@@
Figure 4: Entanglement and non-locality quantifiers for $H\to\gamma\mu\bar{\mu}$ in the $[m_{\mu\bar{\mu}},\cos(\theta_{\gamma\mu})]$ plane, corresponding to the full (first row) and hybrid (second row) computations. Dashed lines are contours corresponding to 0.1 and 0.9 for concurrences and to 2 and 3.5 for Mermin operator.
@@ 5.2 Muon case
Figure 5: Entanglement and non-locality quantifiers for $H\to\gamma e\bar{e}$ in the $[m_{e\bar{e}},\cos(\theta_{\gamma e})]$ plane corresponding to the full computation. Dashed lines are contours corresponding to 0.1 and 0.9 for concurrences and to 2 and 3.5 for Mermin operator.
@@ 6 Summary and perspectives
Our goal was to identify regions of the phase-space where the final particles result entangled after the Higgs boson decay and to determine the feasibility of testing Bell inequality under these kinematical configurations. By expanding beyond traditional bipartite systems to the three-body final state, we compute various entanglement measures including one-to-other and one-to-one concurrences, the area of the concurrence triangle, the conditions for genuine entanglement of 3-qubit systems using the concurrence vector and the three-tangle measure. These quantifiers were derived analytically in terms of the generic 1-loop form factors and numerical predictions in the $[m_{l\bar{l}},\cos(\theta_{\gamma l})]$ plane for three kind of computations were also explored.
@@ Appendix A Conventions and kinematics
In particular, the spinors have eigenvalues ±$\pm 1/2$ and we follow the Chanowitz convention
@@
$p_{H}=\left(\sqrt{m_{\rm H}^{2}+|\vec{k}|^{2}},|\vec{k}|\sin(\theta_{\gamma l}),0,|\vec{k}|\cos(\theta_{\gamma l})\right)\,,\quad k=\left(|\vec{k}|,|\vec{k}|\sin(\theta_{\gamma l}),0,|\vec{k}|\cos(\theta_{\gamma l})\right)\,,$
$p_{-}=\left(\sqrt{s}/2,0,0,\sqrt{s}\beta_{l}/2\right)\,,\quad p_{+}=\left(\sqrt{s}/2,0,0,-\sqrt{s}\beta_{l}/2\right)\,\,\,\text{with}\,\,\,|\vec{k}|=\frac{m_{\rm H}^{2}-s}{2\sqrt{s}}\,,$ (A.4)
@@ Appendix B Helicity amplitudes
[…]
$\mathcal{M}_{++-}|_{\theta_{\gamma l}=0}=-\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
$\mathcal{M}_{--+}|_{\theta_{\gamma l}=0}=\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{-i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
$\mathcal{M}_{+-+}|_{\theta_{\gamma l}=\pi}=-\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
$\mathcal{M}_{-+-}|_{\theta_{\gamma l}=\pi}=\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{-i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
[…]
@@ Appendix C Main results using Dalitz plots representation
Hence Fig. 7 presents the results of the full computation in the $[\sqrt{u},\sqrt{t}]$ plane, already discussed in Figs. 3-5. This format may be more convenient from the experimental point of view in these three-body decays, however we chose the $[m_{l\bar{l}},\cos(\theta_{\gamma l})]$ plane since is more intuitive for identifying kinematical configurations. Of course the physics behind these phenomena is the same in both cases.
@@
Figure 7: Entanglement and non-locality quantifiers in the $[\sqrt{u},\sqrt{t}]$ plane corresponding to the full computation for tau-lepton (first row), muon (second row) and electron (third row).

### v3

@@ 1 Introduction
In this work, we are interested on entanglement between the helicity degrees of freedom of the particles in the tripartite final state corresponding to the Higgs boson decaying into a photon and a lepton-pair. Previously, entanglement properties of bipartite Higgs decays were analyzed for the 2-qubit final states [13, 14, 15], corresponding to tau-lepton pair and two photons. Higgs boson decaying into massive gauge bosons, i.e. 2-qutrits systems, were studied in [16, 17, 18, 19, 20, 21, 22, 23]. Also, 2-qutrits systems were explored in diboson production at LHC [20, 19, 24, 25] and through vector boson scattering [26].
On the other hand, tripartite entanglement within the HEP context is in development: Ref. [27] explores positronium decaying into three photons, QED scattering processes with spectator particle were analyzed in [28, 29, 30], Ref. [31] studies heavy fermion decaying into three fermions via generic (pseudo)scalar, (pseudo)vector and (pseudo)tensor interactions, Ref. [32] explores three-flavor entanglement in neutrino oscillations, Ref. [33] worked in $t\bar{t}Z$ system at the LHC, and Ref. [34] addresses entanglement among the two spins and total angular momentum in $H\to ZZ,WW$. General properties of multipartite systems were presented in [35, 36]. Furthermore, the tripartite Higgs boson decays considered here can be related to bipartite ones, in order to investigate post-decay entanglement [37] and autodistillation phenomena [38, 39].
@@ 2 Three-body Higgs boson decays
The last sum involves the 8 helicity amplitudes (see Appendix B). The polar angle $\theta_{\gamma l}$ belongs to $0,\pi]$ and the Mandelstam variable $s$ has relevant limits
@@
We are interested in quantum entanglement properties of the spin degrees of freedom corresponding to the final particles in the process $H\to\gamma l\bar{l}$. Since photons have two transverse polarizations and leptons have two helicities along the momentum direction, they correspond to qubits in the Quantum Information language. This 3-qubit system is described by the pure state $|\psi\rangle$, which is expanded using the helicity amplitude basis ${+,}{+,}\{+,-\}\otimes\{$$\otimes\{+,-\}$. Therefore, the associated $\times8$ density matrix $\rho=|\psi\rangle\langle\psi|$ for a 3-qubit system, in terms of the helicity amplitudes of the process, is
@@
$\hskip-8.53581pt|\psi\rangle$ $\simeq$ $\frac{-i}{N}\left(8A_{\rm Tree}\kappa_{\rm CP}^{l}e^{-i\delta_{\rm CP}^{l}}\frac{s}{(m_{\rm H}^{2}-s)\sin(\theta_{\gamma l})}|+++\rangle+a_{2}(m_{\rm H}^{2}-s)\sqrt{s}(1+\cos(\theta_{\gamma l}))|++$-$\rangle\right$. (3.5)
@@
$\hskip 34.1433pt\times\left(c_{\gamma l}^{2}\left(4m_{l}^{2}-s\right)\left(s\left(s-16m_{l}^{2}c_{\rm CP}^{2}\right)+m_{\rm H}^{4}+2m_{\rm H}^{2}s\right)\right$.
@@
Concerning the one-to-one concurrences, they require a $\times4$ diagonalization and analytical expressions result just by neglecting lepton mass terms. In that regime, both photon-to-lepton $\mathcal{C}_{\gamma l}$ and photon-to-antilepton $\mathcal{C}_{\gamma\bar{l}}$ exactly vanish for both tree level and 1-loop contributions (then lepton mass terms in the spinors will be relevant in the numerical analysis) and the lepton-to-antilepton $\mathcal{C}_{l\bar{l}}$ for the tree level is very compact
@@
The previous two classes of entanglement (one-to-other and one-to-one) are inequivalent. In fact, the $i$-to-jkjk entanglement limits the entanglements $i$-$j$ and $i$-$k$ by means of the Coffman-Kundu-Wootters (CKW) monogamy inequality [68]
@@ 4.2 $H\to\gamma Z$ and post-decay entanglement
where $\{k^{\mu},s_{1}\}$ and $\{q^{\nu},\bar{s}_{1}\}$ are the momentum and helicity of the photon and $Z$ boson, $\varepsilon$’s are their polarization vectors and the 3-legs form factor $V_{H\gamma Z}$ accounts for the 1-loop contribution (that only depends on $k$ and $q$ momenta). The precise dependence of this form factor on the momenta is not relevant for the following discussion. In the $\{+,-\}\otimes\{+,0,-\}$ basis, the resulting density matrix is
@@
The $H\to\gamma Z$ and our process of interest $H\to\gamma l\bar{l}$ allow to test the post-decay entanglement [37] and autodistillation phenomena [38, 39]. Concretely, the concurrence $\mathcal{C}_{\rm initial}$ of an ‘initial’ bipartite state $\{a,\bar{a}\}$ is computed. The particle $\bar{a}$ decays along the process $\bar{a}\to b_{1}b_{2}$ and the entangled ‘initial’ state led to the entangled ‘final’ subsystems $\{a,b_{1}\}$ and $\{a,b_{2}\}$ with concurrences $\mathcal{C}_{\rm final\,1}$ and $\mathcal{C}_{\rm final\,2}$, respectively. A priori, different amount of entanglement is expected and the autodistillation phenomena arises when final concurrences after decay are greater than the initial one. In our context, we consider the two-body intermediate Higgs decay into the ‘initial’ state $\gamma Z$ and the subsequent post-decay $Z\to l\bar{l}$. Diagrams type (f), with $Z$ boson as mediator, contribute to the 1-loop amplitude in Eq. (2.3) as
@@ 5 Numerical Results
In this section, we present the distributions of the entanglement and Bell non-locality measures in the $m_{l\bar{l}},\cos(\theta_{\gamma l})]$ plane of the phase-space. We analyze the three lepton families separately since the 1-loop contribution dominates different energy regimes in each case, and also corresponds to separate experimental channels. The numerical results correspond to the SM, i.e. $\kappa_{\rm CP}^{l}=1$ and $\delta_{\rm CP}^{l}=0$ in Eq. (B.1), including lepton mass effects and using the corresponding form factors for each kind of computation in Eq. (2.4). Then, in the plots, the range for $\cos(\theta_{\gamma l})$ is $[0,1]$ since the distributions result symmetric under $\cos\to-\cos$, and the dilepton invariant mass covers the range in Eq. (2.6).
@@
Figure 3: Entanglement and non-locality quantifiers for $H\to\gamma\tau\bar{\tau}$ in the $m_{\tau\bar{\tau}},\cos(\theta_{\gamma\tau})]$ plane, corresponding to the full computation (first row) and two-body intermediate Higgs decay (second row). Dashed lines are contours corresponding to 0.1 and 0.9 for concurrences and to 2 and 3.5 for Mermin operator.
@@ 5.1 Tau-lepton case
The $H\to\gamma\tau\bar{\tau}$ process is dominated by the 1-loop (tree level) contribution for energies below (above) 30 GeV, as can be seen from the first row of Fig. 2. Around the $Z$-pole, both contributions interfere constructively but the resonance comes from the 1-loop diagrams type (f). The resulting one-to-other and one-to-one concurrences $\mathcal{C}_{l\bar{l}|\gamma}$ and $\mathcal{C}_{l\bar{l}}$ in the full computation are shown in the first row of Fig. 3, together with the Mermin operator $\mathcal{M}_{3}$ in the $m_{\tau\bar{\tau}},\cos(\theta_{\gamma\tau})]$ plane.
@@
Figure 4: Entanglement and non-locality quantifiers for $H\to\gamma\mu\bar{\mu}$ in the $m_{\mu\bar{\mu}},\cos(\theta_{\gamma\mu})]$ plane, corresponding to the full (first row) and hybrid (second row) computations. Dashed lines are contours corresponding to 0.1 and 0.9 for concurrences and to 2 and 3.5 for Mermin operator.
@@ 5.2 Muon case
Figure 5: Entanglement and non-locality quantifiers for $H\to\gamma e\bar{e}$ in the $m_{e\bar{e}},\cos(\theta_{\gamma e})]$ plane corresponding to the full computation. Dashed lines are contours corresponding to 0.1 and 0.9 for concurrences and to 2 and 3.5 for Mermin operator.
@@ 6 Summary and perspectives
Our goal was to identify regions of the phase-space where the final particles result entangled after the Higgs boson decay and to determine the feasibility of testing Bell inequality under these kinematical configurations. By expanding beyond traditional bipartite systems to the three-body final state, we compute various entanglement measures including one-to-other and one-to-one concurrences, the area of the concurrence triangle, the conditions for genuine entanglement of 3-qubit systems using the concurrence vector and the three-tangle measure. These quantifiers were derived analytically in terms of the generic 1-loop form factors and numerical predictions in the $m_{l\bar{l}},\cos(\theta_{\gamma l})]$ plane for three kind of computations were also explored.
@@ Appendix A Conventions and kinematics
In particular, the spinors have eigenvalues $\pm 1/2$ and we follow the Chanowitz convention
@@
$\hskip-14.22636ptp_{H}=\left(\sqrt{m_{\rm H}^{2}+|\vec{k}|^{2}},|\vec{k}|\sin(\theta_{\gamma l}),0,|\vec{k}|\cos(\theta_{\gamma l})\right)\,,\quad k=\left(|\vec{k}|,|\vec{k}|\sin(\theta_{\gamma l}),0,|\vec{k}|\cos(\theta_{\gamma l})\right)\,$,
$\hskip-14.22636ptp_{-}=\left(\sqrt{s}/2,0,0,\sqrt{s}\beta_{l}/2\right)\,,\quad p_{+}=\left(\sqrt{s}/2,0,0,-\sqrt{s}\beta_{l}/2\right)\,\,\,\text{with}\,\,\,|\vec{k}|=\frac{m_{\rm H}^{2}-s}{2\sqrt{s}}\,$, (A.4)
@@ Appendix B Helicity amplitudes
Some comments about interesting limits are in order. First, when photon has vanishing energy ($s=m_{\rm H}^{2}$), the 1-loop contribution vanishes for all helicity amplitudes and the tree level yields to IR divergences for ${+++,+,+++,}\{$+++,+--,-+++,---\} amplitudes (which are avoided by the lower cut $E_{cut}^{\gamma}$). Secondly, when photon is collinear with lepton ($\theta_{\gamma l}=0$) or with antilepton ($\theta_{\gamma l}=\pi$), just two helicity amplitudes are non-vanishing for each case:
$\hskip-22.76219pt\mathcal{M}_{++-}|_{\theta_{\gamma l}=0}=-\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
$\hskip-22.76219pt\mathcal{M}_{--+}|_{\theta_{\gamma l}=0}=\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{-i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
$\hskip-22.76219pt\mathcal{M}_{+-+}|_{\theta_{\gamma l}=\pi}=-\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
$\hskip-22.76219pt\mathcal{M}_{-+-}|_{\theta_{\gamma l}=\pi}=\frac{i\sqrt{2}A_{\rm Tree}\kappa_{\rm CP}^{l}e^{-i\delta_{\rm CP}^{l}}\sqrt{s}}{m_{l}}$
@@
lepton case $\sqrt{s}_{\rm min}$ [GeV] $\sqrt{s}_{\rm max}$ [GeV] $\Gamma_{\rm full}$ [KeV] $\Gamma_{\rm hybrid}$ [KeV] $N_{\rm Run\,2+3}$
@@ Appendix C Main results using Dalitz plots representation
Hence Fig. 7 presents the results of the full computation in the $\sqrt{u},\sqrt{t}]$ plane, already discussed in Figs. 3-5. This format may be more convenient from the experimental point of view in these three-body decays, however we chose the $m_{l\bar{l}},\cos(\theta_{\gamma l})]$ plane since is more intuitive for identifying kinematical configurations. Of course the physics behind these phenomena is the same in both cases.
@@
Figure 7: Entanglement and non-locality quantifiers in the $\sqrt{u},\sqrt{t}]$ plane corresponding to the full computation for tau-lepton (first row), muon (second row) and electron (third row).